Good Sports, Inc., is a private full-line sporting goods retailer. Assume one of the Good Sports stores reported current asseCan someone please help? I can't figure out Transaction (4) and it is not 1.07 nor is it 1.03.

Good Sports, Inc., is a private full-line sporting goods retailer. Assume one of the Good Sports stores reported current assets of $87,850 and its current ratio was 1.75, and then completed the following transactions: (1) paid $6,100 on accounts payable, (2) purchased a delivery truck for $11,000 cash, (3) wrote off a bad account receivable for $1,800, and (4) paid previously declared dividends in the amount of $23,500. Required: Compute the updated current ratio after each transaction, by showing the cumulative effects of the transactions in the following table. (Round your answers to 2 decimal places.)
Answer #1

Current liabilities = 87850/1.75 = 50200

calculate current ratio

Current ratio
Transaction 1 (81750/44100) = 1.85
Transaction 2 (70750/44100) = 1.60
Transaction 3 70750/44100 = 1.60
Transaction 4 (47250/20600) = 2.29